# Accommodating Fabrication Defects on Floquet Codes with Minimal Hardware Requirements

### v1

@@ Appendix G Strategy for removing defective connections
In the surface code, a defective connection between a data qubit and a measurement qubit can be removed by treating the data qubit as defective [19]. One can apply the same technique to Floquet codes (see Sec. 5.2), but improvements on this strategy are possible. With an improved strategy, one can remove defective connections from a Floquet code with no additional connections, no schedule modifications, and without removing any data qubits.
@@
Having chosen our merge and split plaquettes, we then proceed using Steps 4-7 of the algorithm in Sec. 3.1. First, in Fig. 21(b), we split the blue plaquette into weight-2 plaquettes by introducing additional red edges. Then, in Fig. 21(c) we use the newly-added red edges to merge the green plaquettes into a super-plaquette. The resulting graph is tri-valent and three-face-colourable, meaning that a valid Floquet code can be constructed from it, without needing to remove any data qubits.
The result has some similarities with the strategy for removing defective qubits outlined in Sec. 3. First, any newly-added edges are between qubits which are already connected, so no additional hardware connections or measurement schedule modifications are required. We can also see that this strategy can be run iteratively to accommodate multiple defective connections. Finally, this strategy can be applied not just to the honeycomb code, but to other Floquet codes defined on the 2-D colour code lattice.

### v2

@@ Appendix G Strategy for removing defective connections
In the surface code, a defective connection between a data qubit and a measurement qubit can be removed by treating the data qubit as defective [20]. One can apply the same technique to Floquet codes (see Sec. 5.2), but improvements on this strategy are possible. With an improved strategy, one can remove defective connections from a Floquet code with no additional connections, no schedule modifications, and without removing any data qubits. We will focus on removing a single defective connection in this example, but note that as before this approach can be used iteratively to accommodate multiple defective connections. We also note that this approach can be used in combination with the strategy in Sec. 3 to correct for a combination of defective qubits and connections. Finally, while this example will focus on the honeycomb code, we note that it can also be applied to Floquet codes defined on other tri-valent and 3-face-colourable lattices.
@@
Having chosen our merge and split plaquettes, we then proceed using Steps 4-7 of the algorithm in Sec. 3.1. First, in Fig. 21(b), we split the blue plaquette into weight-2 plaquettes by introducing additional red edges, akin to Steps 4 and 5. Then, in Fig. 21(c) we use the newly-added red edges to merge the green plaquettes into a super-plaquette, using the same process as in Steps 6 and 7. The theorem below is equivalent to Theorem 3, but looking at standard colour code lattices rather than the “heavy” lattices considered in Sec. 5.2. This difference does not affect the proof.
Theorem 3 (for defective connections).
The above process produces a tri-valent 3-face-colourable lattice.
Proof.
Our proof will follow a similar structure to the proof for Theorem 1. We will prove this for removing a single defective edge, whose basis is denoted as $D$.
Note that edges removed from the lattice are edges which contribute to the shrink plaquette and whose basis is $D$. The shrink plaquette consists of a cycle of edges whose bases alternate between $D$ and one other basis $T$. This means that each vertex in the shrink plaquette is incident to one edge which is removed from the lattice. However, each vertex in the shrink plaquette is also incident to a newly-added edge of type $D$, therefore vertices in the shrink plaquette are still of degree three. Vertices wich are not in the shrink plaquette are not affected by this, therefore tri-valence is still preserved.
3-face-colourability follows the same proof as in Theorem 1. The newly-added weight-2 plaquettes are assigned the same colour as the shrink plaquette, and the super-plaquette is assigned the same colour as the merge plaquette. Because the merge plaquettes are adjacent to the shrink plaquette, they must be different colours. The weight-2 plaquettes are adjacent to the super-plaquette and one plaquette which was adjacent to the shrink plaquette in the original graph, both of which must be coloured differently because of the remarks above. The super-plaquette on the other hand is adjacent to weight-2 plaquettes—which are of the same colour as the shrink plaquette—and plaquettes which were incident to a merge plaquette in the original graph, and must therefore also be a different colour. This means that the newly-added plaquettes are guaranteed to not share the same colour as any adjacent plaquette, thus preserving 3-face-colourability. ∎
Note that unlike the strategy for qubit removal, the total number of vertices, edges and plaquettes remains unchanged compared to the original lattice when removing a defective edge. We can also see that our scheme is optimal in the number of edges removed from the lattice, again considering honeycomb, 4.8.8, or 4.6.12 lattices. This is a rephrasing of Theorem 4, with a focus on the standard honeycomb, 4.8.8 and 4.6.12 lattices rather than their heavy equivalents. This modification does not affect the proof.
Theorem 4 (for defective connections).
Consider a defective edge in the bulk of a Floquet code defined on a honeycomb, 4.8.8 or 4.6.12 lattice (without boundary); i.e., a uniform tiling of the plane. Under the constraint that we cannot form an edge between any two qubits not connected by an edge in the original lattice, the algorithm described above removes the defective edge with the minimal number of removed edges from the original lattice.
Proof.
Our proof will follow a similar form to that of Theorem 2. The defective edge $E$ is incident to two vertices $v_{1},v_{2}$. Removing $E$ will result in $v_{1}$ and $v_{2}$ both being of degree-2. In order to maintain tri-valence we must introduce additional edges incident to $v_{1}$ and $v_{2}$. Again, we assume the edges in the original lattice, with the exception of the defective edge, account for all possible available connections in hardware, and any new edges must lie along the same connections.
Without loss of generality, let $E$ be coloured red, as in Fig. 21. With the defective edge removed, $v_{1}$ and $v_{2}$ are not incident to any red edge, meaning that red edges must be added to the graph. To maintain tri-valence and three-face-colourability, we must do this by drawing a path of edges between $v_{1}$ and $v_{2}$ which (a) alternates between red and non-red edges, (b) begins and ends with non-red edges, and (c) does not traverse $E$. Using this path, a new tri-valent graph can be constructed by removing red edges along the path and introducing new red edges parallel to the non-red edges along the path. The number of edges removed from the original lattice will therefore be $(n-2)/2$, where $n$ is the number of vertices in the path. Therefore choosing the shortest path will minimise the number of edges removed.
In the approach described above, we construct this path out of the shrink plaquette. $v_{1}$ and $v_{2}$ are incident to an edge which is part of the shrink plaquette, meaning that the length of the path will be the weight of the shrink plaquette. Therefore the number of removed edges along the path will be $(K-2)/2$, where $K$ is the weight of the shrink plaquette. This leads to 2 additional edges being removed in the honeycomb lattice, as all plaquettes are of weight 6. This path has length 5. A shorter path would have to have length 3, implying the existence of a length-4 cycle in the original graph (since $v_{1}$ and $v_{2}$ are connected by an edge in the original graph). This does not exist in the honeycomb lattice. In the 4.8.8 lattice, the number of additional edges removed will either be 1 if the defective edge is part of a weight-4 plaquette, or 3 otherwise. In the latter case, there is no shorter path between $v_{1}$ and $v_{2}$ other than along an arbitrarily chosen shrink plaquette (both having weight 8). Finally, for the 4.6.12 lattice, the number of additional edges removed will either by 1 if the defective edge is part of a weight-4 plaquette, or 2 if the edge is incident to a weight-6 plaquetteNote that the structure of the 4.6.12 lattice ensures that any edge is incident to either a weight-4 or weight-6 plaquette, so we do not have to consider choosing a weight-12 plaquette as the shrink plaquette.. In this case as well, it is obvious that a shorter path which connects $v_{1}$ and $v_{2}$ while meeting the requirements above cannot exist. ∎
It can also be easily seen that this optimality is not constrained to planar surfaces. Below, we show a simple remark found by applying this proof to the 8.8.8 lattice used in hyperbolic Floquet codes [48, 49]:
Remark 2.
Consider a defective edge in the bulk of a Floquet code defined on an 8.8.8 lattice (without boundary), which is a particular uniform tiling of a hyperbolic surface. Under the constraint that we cannot form an edge between any two qubits not connected by an edge in the original lattice, the algorithm described above removes the defective edge with the minimal number of removed edges from the original lattice.
Proof.
Our proof will follow from the proof of Theorem 4, noting that the proof does not rely on the lattice being a plane. In the 8.8.8 lattice, the shrink plaquette will always be of weight-8, meaning that there will be 8 vertices in the path connecting the vertices incident to the defective edge, and a total of 3 edges are removed from the lattice. It is impossible to construct a shorter path in the 8.8.8 lattice satisfying the requirements stated in Theorem 4, therefore this edge removal is optimal. ∎
Note that we do not claim optimality for the scenario for non-trivial combinations of defective connections. As with Theorem 2, an optimal edge removal in the case of multiple defective edges would depend on the choice of shrink and merge plaquettes. In particular, if the shrink plaquette is of large weight then there might be a shorter path between the two vertices incident to the defective edge. In general, for $n$ defective connections the number of possible sets of shrink plaquettes grows as $O(2^{n})$, so it is likely that one will need to rely on heuristics to make suitable choices.